### v2

# AfroXLMR-Social: Adapting Pre-trained Language Models for African Languages Social Media Text

AfroXLMR-Social: Adapting Pre-trained Language Models
for African Languages Social Media Text
Tadesse Destaw Belay1, Israel Abebe Azime2, Ibrahim Said Ahmad3,4, David Ifeoluwa Adelani5,6
Idris Abdulmumin7
, Abinew Ali Ayele8, Shamsuddeen Hassan Muhammad4,9,
Seid Muhie Yimam10
1Instituto Politécnico Nacional, 2Saarland University, 3Northeastern University, 4Bayero University Kano, 5Mila, McGill University,
6Canada CIFAR AI Chair, 7University of Pretoria, 8Bahir Dar University, 9Imperial College London, 10University of Hamburg,
Contact: tadesseit@gmail.com
AfroXLMR-Social: Adapting Pre-trained Language Models
for African Languages Social Media Text
Tadesse Destaw Belay1, Israel Abebe Azime2, Ibrahim Said Ahmad3,4, David Ifeoluwa Adelani5,6 Idris Abdulmumin7, Abinew Ali Ayele8, Shamsuddeen Hassan Muhammad4,9, Seid Muhie Yimam10 1Instituto Politécnico Nacional, 2Saarland University, 3Northeastern University, 4Bayero University Kano, 5Mila, McGill University, 6Canada CIFAR AI Chair, 7University of Pretoria, 8Bahir Dar University, 9Imperial College London, 10University of Hamburg, Contact: tadesseit@gmail.com
Language models built from various sources are the foundation of today’s NLP progress. However, for many low-resource languages, the diversity of domains is often limited—more biased to a religious domain, which impacts their performance when evaluated on distant and rapidly evolving domains such as social media. Domain adaptive pre-training (DAPT) and task-adaptive pre-training (TAPT) are popular techniques to reduce this bias through continual pre-training for BERT-based models, but they have not been explored for African multilingual encoders. In this paper, we explore DAPT and TAPT continual pertaining approaches for the African languages social media domain. We introduce AfriSocial—a large-scale social media and news domain corpus for continual pre-training on several African languages. Leveraging AfriSocial, we show that DAPT consistently improves performance on three subjective tasks: sentiment analysis, multi-label emotion, and hate speech classification, covering 19 languages from 1% to 30% F1 score. Similarly, leveraging TAPT on one task data improves performance on other related tasks. For example, training with unlabeled sentiment data (source) for a fine-grained emotion classification task (target) improves the baseline results by an F1 score ranging from 0.55% to 15.11%. Combining these two methods (i.e. DAPT + TAPT) further improves the overall performance.
1 Introduction
Pre-trained language models (PLMs) are initially trained on vast and diverse corpora, including encyclopedias and web content Conneau et al. (2020a); Chiang et al. (2022). Subsequently, these pre-trained models are used in supervised training for a specific Natural Language Processing (NLP) task by further finetuning. Fine-tuned PLMs achieve strong performance across many tasks and datasets from various sources Wu et al. (2024). However, this raises a question: Do PLMs function universally across domains, or does continual training of PLMs with domain-specific data offer better performance?
Figure 1: Continual pretraining illustration. A general-purpose pretrained model, such as AfroXLMR, is first adapted to a social domain, resulting in AfroXLMR-Social. This model then undergoes Cross-task TAPT using sentiment analysis, emotion, and hate speech data without the labels for further fine-tuning.
While some studies have shown the benefit of continual pretraining on a domain-specific unlabeled data Lee et al. (2019); Gururangan et al. (2020), one domain may not be generalizable to other domains and languages. Moreover, it is unknown how the benefit of continual pretraining may vary with factors like the amount of unlabeled corpus, the source domain itself, the evaluation task, the resource richness of the target languages, and the trained target model Gururangan et al. (2020). This raises the question of whether pretraining on a corpus more directly tied to the task can further improve performance. This work addresses these questions by continual domain adaptive pertaining (DAPT) and task adaptive pretraining (TAPT) on the downstream subjective NLP tasks in a low-resource language setup. We consider the social media domain X and News for a continual pretraining from a high-performing multilingual baseline model, AfroXLMR Alabi et al. (2022), a continual pretrained model for African languages from XLM-RoBERTa Conneau et al. (2020a). We explore subjective NLP task results from baseline (the base model result), DAPT, TAPT, and DAPT + TAPT, on a smaller but directly domain- and task-relevant unlabeled corpus. Our result shows that DAPT and TAPT highly benefit from similar source NLP tasks. Figure 1 shows the general high-level continual pre-training illustration. In summary, our contributions are:
  • Presenting AfriSocial, a new quality domain-specific corpus for 14 African languages, collected from the social domain (X) and news.
  • A through analysis of domain and task adaptive continual pretraining across subjective NLP tasks for low-resource languages.
  • Achieving state-of-the-art results in the evaluated NLP tasks and publicly making AfriSocial social-domain corpus and AfroXLMR-Social pretrained models for developing low-resource languages.
2 Related Work
Domain in NLP Language models (LMs) pretrained on text from various sources are the foundation of today’s NLP. Domain adaptation in NLP refers to enhancing the performance of a model using similar domain data (target domain) by leveraging knowledge from an existing domain (source domain) Ramponi and Plank (2020). Domain refers to different implicit clusters of text representations in pretrained LMs, such as news articles, social media, medical texts, or legal documents Aharoni and Goldberg (2020); Shi et al. (2024). Each domain has its unique characteristics, vocabulary, and writing style, which can affect the performance of NLP models when applied to new or unseen domains. Therefore, a similar domain means the text source from which the pre-trained model was made, similar to the target NLP task data source.
Continual Pre-training of LMs There are several techniques for the downstream NLP task improvements, such as general-purpose pre-training Wang et al. (2023), language-adaptive continual pre-training Alabi et al. (2022), domain-adaptive pre-training Gururangan et al. (2020), task-adaptive pre-training Alabi et al. (2022), and data augmentation Zhang et al. (2024). Prior works have shown the benefit of continual pretraining using a domain-specific corpus Gururangan et al. (2020) and training LMs in a specific domain from scratch Huang et al. (2020). However, continual pretraining is arguably more cost-effective since it is a continuous pretraining from the existing base language model.
Domain Adaptive Pre-training (DAPT) Domain Adaptive pre-training (DAPT) is straightforward, continuing pretraining a model on a corpus of unlabeled domain-specific text Aharoni and Goldberg (2020). DAPT techniques handle discrepancies between the source (pre-training) and target (fine-tuning) domains. Traditional fine-tuning often yields suboptimal results when pre-trained models encounter data that diverges significantly from their training data. DAPT techniques reduce this mismatch by aligning data distributions, ensuring the model can generalize better in a novel setting. Lee et al. (2019) considers a single domain at a time and uses a language model pretrained on a smaller and less diverse corpus than the most varied and multilingual language models.
Task Adaptive Pre-training (TAPT) Task-adaptive pretraining (TAPT) refers to pretraining on the unlabeled training set for a similar task-specific data Gu et al. (2024). Compared to DAPT, TAPT solely leverages the training data of the similar downstream task for continuous pre-training. It uses a far smaller pretraining corpus that is much more task-relevant (assuming that the training set represents aspects of the task well). This makes TAPT much less expensive to run than DAPT. There are also a combination of various techniques, such as DAPT followed by TAPT, which is beneficial for end-task performance Gururangan et al. (2020).
Language Adaptive Pre-training (LAPT) LAPT is also called language-adaptive fine-tuning (LAFT) Alabi et al. (2022). It mainly focuses on adapting a pre-trained language model to a specific language(s) using any language-specific corpus. This is the most common approach: collecting any corpus for fine-tuning language models without considering the domain Yu and Joty (2021); Alabi et al. (2022). LAPT is a vital step to improve language understanding and representation, especially for low-resource languages Wang et al. (2023). However, in the primary studies that investigated LAPT with various language-specific corpora, the impacts of the text sources on specific NLP tasks are unexplored.
Although XLMR Conneau et al. (2020a) pretraining corpus is derived from multiple sources and languages, it has not yet been explored whether these sources are diverse enough to generalize in a specific domain and task. This leads to asking whether subjective tasks related to social media text can be understood with this generic model. Towards this end, we explore further a continual training of DATP and TAPT from AfroXLMR Alabi et al. (2022) and evaluating the impacts on highly subjective NLP tasks such as sentiment analysis, multi-label emotion, and hate speech classification.
3 AfriSocial: Domain Adaptation Corpus
Social Domain in our Work
In computational linguistics, domain boundaries can be defined through various dimensions including content, style, and purpose Plank (2016). In this work, we define the social domain based on two key factors: 1) Convergent textual characteristics, X and news sources provide a public discourse, comments, reactions, and conversational linguistic patterns that facilitate social interactions, and cultural expressions and 2) Functional similarity in downstream tasks, the selected evaluation task datasets are sourced from the two sources, shown in Table 1. This grouping is empirically supported by Derczynski et al. (2016), which demonstrated comparable performance patterns when evaluating entity recognition models across X and news data, suggesting underlying linguistic commonalities despite surface differences.
We create AfriSocial, a social domain-specific corpus comprising X and news for 14 African languages, shown in Table 2. We select X and news because they are the most common text sources for low-resource languages to annotate a dataset for supervised NLP tasks. The motivations behind creating this domain-specific corpus are the following.
Limited coverage for African languages
The available well-known compiled corpora are limited to include African languages; most of them are only English-centric, such as fineweb Penedo et al. (2024) and C4 Raffel et al. (2020). The reasons include the extra effort required to collect data for such low-resource languages, the limited availability of text, and the challenges in detecting the language and filtering sources.
Text quality issues
The quality of the available corpus is under consideration, especially for low-resource languages. For example, in the OPUS corpora Lison and Tiedemann (2016), there are Tigrinya (tir) texts under the Amharic (amh) file as both languages use the same script. Some of the available corpora are translated, such as OPUS-100 Zhang et al. (2020), with the problem that the quality of the translator tool is still not mature enough for low-resource languages.
Non availability of social domain corpus
To be specific, a domain-specific corpus is vital for adapting pre-trained language models into a specific domain, such as a health-specific domain. Likewise, the social media corpus is limited even to a highly resourced language.
Dataset name Task name # lang. Data Sources # Classes
AfriSenti Muhammad et al. (2023); Yimam et al. (2020) Sentiment analysis 14 News, social media 3
AfriEmo Muhammad et al. (2025b); Belay et al. (2025a) Emotion analysis 17 News, social media 6
AfriHate Muhammad et al. (2025a) Hate speech detection 15 News, social media 3
AfriSocial (new domain specific corpus) Unlabeled 14 X and news -
Table 1: List of datasets. Social media sources include posts/comments from YouTube and X. The AfriSenti dataset class labels are positive, negative, and neutral. The AfriEmo dataset labels are six basic emotions (anger, disgust, fear, joy, sadness, and surprise) in a multi-label annotation. AfriHate labels are abuse, hate, and neutral.
3.1 Data Sources Selection
Based on our assessment, most of the NLP datasets of African languages are sourced and annotated from X and news domain, as sufficient text can be found in these two sources. As shown in Table 1, subjective NLP tasks for African languages, such as sentiment analysis, multi-label emotion, and hate speech classification datasets, are sourced from X and news. The AfriSocial corpus is sourced from similar domains to improve these subjective tasks further, X and News. There is an X domain corpus and model for high-resource languages such as XLM-T Barbieri et al. (2022) to evaluate and improve task datasets sourced from X. However, no available corpora specialize in the social domain for low-resource African languages. More details about the data collections are presented in Appendix I.
Lang. X News Total Sent.
amh 588,154 45,480 633,634
ary 9,219 156,494 165,712
hau 640,737 30,935 671,672
ibo 15,436 38,231 53,667
kin 16,928 72,583 89,511
orm 33,587 59,429 93,016
pcm 106,577 7,781 116,358
som 144,862 24,473 169,335
swa 46,588 - 46,834
tir 167,139 45,033 212,172
twi 8,681 - 8,681
yor 26,560 49,591 76,151
xho - 354,959 354,959
zul 12,102 854,587 866,689
Total 1.82M 1.74M 3.56M
Table 2: AfriSocial corpus statistics at language and source level, where Total Sent. is the number of sentences. The full names of the languages are presented in Appendix A.
3.2 Pre-processing and Quality Control
We apply the following quality measures on the AfriSocial corpus.
Language Identification (LID)
We apply better LID tools for each language. For example, for the language mixing problem of the existing corpus mentioned in Sec §3, we used separate LID tools, GeezSwitch [^1] https://pypi.org/project/geezswitch/ to handle Ethiopic script languages (Amharic and Tigrinya) and pycld3222https://pypi.org/project/pycld3/ for the supported Latin and other script languages at the sentence level.
Sentence Segmentation
The same approach as language identification, we used proper tools for each language to segment into sentences. For the Ethiopic script language, we used amseg tool Yimam et al. (2021), and for other Latin script languages, we used NLTK Bird and Loper (2004).
Other Preprocessing
We exclude short sentences, only URL lines, and anonymize personally identifiable information (PII) such as usernames starting with the @ symbol, and email addresses. We pay special attention to ensuring that the available evaluation task data (sentiment, emotion, and hate speech) do not appear in the AfriSocial corpus before and after processing. De-duplication is applied if a near-similar instance is present, excluding it from AfriSocial, not from the annotated dataset. Table 2 shows the AfriSocial corpus statistics with their sources and number of sentences.
4 Evaluation Tasks and Datasets
We select subjective NLP tasks for our evaluation based on the following reasons. 1) Subjective tasks face more disagreement during annotations, leading to less performance in the evaluation, especially for low-resource languages Fleisig et al. (2023). As we can see from the SemEval-2025 Task 11 Muhammad et al. (2025c), an emotion detection shared task covering 32 languages, low-resource languages are not well explored, and the lowest results are from African languages. 2) African language subjective tasks are sourced from X and news, as shown in Table 1. These sources align with the same domain as the AfriSocial corpus. The three subjective tasks for our evaluation are sentiment analysis, multi-label emotion, and hate speech classification. We keep the original train-test split of all evaluation datasets throughout our experiment for later comparison with benchmark results.
4.1 AfriSenti: Sentiment Analysis Dataset
AfriSenti Muhammad et al. (2023) is a sentiment analysis dataset across 14 African languages. The data is sourced from X (formerly Twitter) and annotated in one of the three sentiment classes: positive, negative, and neutral. From 14 languages, the two languages, Oromo (orm) and Tigrinya (tir) have only test sets.
4.2 AfriEmo: Multi-label Emotion Dataset
SemEval-2025-Task 11 Muhammad et al. (2025c) is an emotion dataset that covers 32 languages, from diverse domains such as social media platforms (X, Reddit, YouTube, and others) and news. The AfriSocial domain-specific corpus targets low-resource African languages; we target the African languages emotion dataset from the SemEval-2025 Task 11, which we call AfriEmo. It covers 17 African languages from the 32 languages. This dataset is annotated in a multi-label approach - a text might have any combination (none, one, some, or all) of emotion labels from a given set of emotions (anger, disgust, fear, joy, sadness, and surprise).
4.3 AfriHate: Hate Speech Dataset
AfriHate Muhammad et al. (2025a) is a multilingual hate and abusive speech dataset in 15 African languages sourced from X. Each text is categorized into one of the abusive, hate, or neutral labels. The languages covered in the corresponding evaluation datasets, such as language name, ISO code, countries/regions spoken, language family, and writing script, are presented in the Appendix (A, B).
5 Language Models
5.1 Encoder-only Language Models
Multilingual encoder-only pretrained language models (PLMs) such as XLM-R Conneau et al. (2020a) and mBERT Devlin et al. (2019) have shown impressive capability on many languages for a variety of downstream tasks. They are also often used to initialize checkpoints to adapt to other languages, such as AfroXLMR Alabi et al. (2022), which is initialized from XLM-R to specialize in African languages. We evaluate popular multilingual and African-centric PLMs such as AfroLM Dossou et al. (2022) and AfriBERTa Ogueji et al. (2021) and found that AfroXLMR is better for the targeted evaluation datasets as it covers more African languages. We make AfroXLMR our benchmarks and further train on the AfriSocial domain and the selected evaluation tasks. The AfroXLMR plus DAPT with the AfriSocial domain-specific corpus gives us AfroXLMR-Social. The detailed hyperparameters of the continual training are shown in Appendix D.
5.2 Large Language Models
We compare our DAPT and TAPT approach results from AfroXLMR with state-of-the-art open source LLMs such as Llama 3 Grattafiori et al. (2024), Gemma 2 Team et al. (2024), Mistral Jiang et al. (2024), and proprietary LLMs such as Gemini 1.5 Georgiev et al. (2024) and GPT-4o Hurst et al. (2024). For AfriSenti and AfriHate task results, we use the LLMs benchmark results from the AfroBench Ojo et al. (2025) leaderboard. For the AfriEmo task, we use LLM results from the SemEval-2025 Task 11 datasets papers Muhammad et al. (2025b); Belay et al. (2025a). The detailed versions of the LLMs are presented in the Appendix E.
6 Experiment Results
6.1 Domain Adaptive Pre-training (DAPT)
The domain-adaptive pretraining (DAPT) approach is straightforward; we continue pretraining from the strong baseline language model (AfroXLMR) using the domain-specific AfriSocial corpus in a multilingual setup. Baseline results from (AfroXLMR) and after applying DAPT results are presented in Table 3.
AfriSenti AfriEmo AfriHate
Lang. AfroXLMR DAPT Lang. AfroXLMR DAPT Lang. AfroXLMR DAPT
amh 50.09 57.22 afr 43.66 44.57 amh 73.54 78.57
arq 52.22 64.62 amh 68.97 71.67 arq 43.41 45.96
ary 52.86 62.34 ary 47.62 52.63 ary 75.13 75.6
hau 79.34 81.66 hau 64.30 70.74 hau 81.55 80.78
ibo 76.92 79.8 ibo 26.27 54.54 ibo 82.78 88.05
kin 70.95 72.73 kin 52.39 56.73 kin 75.28 78.75
pcm 50.47 52.09 orm 52.28 61.38 orm 67.23 74.11
por 60.93 64.81 pcm 55.39 59.93 pcm 64.85 67.61
swa 28.26 61.42 ptMZ 22.09 36.80 som 55.66 55.64
tso 35.37 38.81 som 48.78 54.86 swa 91.51 91.2
twi 47.2 56.00 swa 30.74 34.35 tir 50.2 55.9
yor 72.27 74.63 tir 57.22 60.71 twi 46.89 48.42
orm 20.09 24.28 vmw 21.18 22.08 xho 50.91 59.17
tir 22.45 24.53 yor 28.65 39.26 yor 53.44 77.9
Avg. 51.39 58.21 Avg. 44.25 51.45 Avg. 65.17 69.83
Table 3: Result of baseline (AfroXLMR) and DAPT (AfroXLMR-Social) across the three datasets (AfriSenti, AfriEmo, and AfriHate). During TAPT, the text for the task-adaptive data is without the labels, and the evaluation is cross-tasked among the three target datasets. Reported results are macro-F1.
Baseline: As a baseline, we use the results from AfroXLMR Alabi et al. (2022), which perform better than other BERT-based encode-only models across targeted datasets since it includes more African languages during pre-training.
Results: The results before and after DAPT are shown in Table 3, AfroXLMR and DAPT columns, respectively. We observe that DAPT improves over the baseline in all languages and datasets, demonstrating the benefit of DAPT when the target domain is relevant. It consistently improves over the baseline models for each language. It suggests that continual pretraining on a small, quality, domain-relevant dataset is important for subjective tasks from the same domain.
6.2 Task Adaptive Pretraining (TAPT)
Similar to DAPT, TAPT consists of a second phase of continual pertaining. TAPT is a cross-task transfer across datasets, which refers to finetuning on the unlabeled data of the non-targeted task during evaluation. We explore the TAPT approach by directly applying it to the base model and after DAPT. For example, if we make TAPT for the AfriSenti task, we further fine-tune the base model and DAPT model using the unlabeled data of the AfriEmo and AfriHate datasets separately. Compared to DAPT, the task-adaptive approach strikes a different trade-off: it uses a far smaller pretraining corpus, assuming the training set represents aspects of the target task. This makes TAPT much less expensive to run than DAPT.
AfriSenti TAPT AfriEmo TAPT AfriHate TAPT
Lang. Base AfriEmo AfriHate DATP + TAPT Base AfriSenti AfriHate DATP + TAPT Base AfriSenti AfriEmo DATP + TAPT
amh 50.09 54.91 54.48 55.80 68.97 69.56 67.21 71.04 73.54 73.13 73.26 78.06
arq 52.22 62.42 59.41 63.38 44.93 51.25 48.95 48.72 43.41 46.19 43.84 44.21
ary 52.86 64.13 52.80 63.05 47.62 50.21 48.81 51.63 75.13 75.13 71.70 77.51
hau 79.34 80.65 80.08 82.71 64.30 66.93 61.16 69.77 81.55 77.76 81.94 82.09
ibo 76.92 80.01 78.29 80.42 26.27 53.13 51.26 54.26 82.78 87.69 86.52 87.68
kin 70.95 71.47 69.72 69.53 52.39 53.27 53.49 54.47 75.28 77.48 76.30 78.36
orm 20.09 23.53 42.99 28.93 52.28 56.43 52.22 58.75 77.08 69.03 67.67 71.07
pcm 50.47 50.97 50.29 52.04 55.39 56.71 56.60 58.89 64.85 67.73 66.94 69.91
ptMZ 60.93 64.05 62.80 63.75 22.09 37.20 30.99 37.76 - - - -
som - - - - 48.78 50.33 49.63 52.65 55.66 57.29 54.97 56.75
swa 28.26 59.33 57.26 54.94 30.74 33.02 31.85 32.74 91.51 91.91 91.27 91.16
tir 22.45 10.81 16.22 28.90 57.22 55.72 55.84 57.12 50.20 54.21 56.98 32.70
twi 47.20 47.68 50.23 54.47 - - - - 46.89 49.30 48.94 49.01
yor 72.27 72.22 70.90 73.65 28.65 34.34 32.87 35.89 53.44 53.69 54.51 54.76
Avg. 52.62 57.09 57.34 59.35 45.74 51.39 49.22 52.87 66.72 67.46 67.14 67.77
Table 4: Cross-TAPT results across the three datasets. Base column is baseline results from AfroXLMR, DATPT + TAPT column results for AfriSenti and AfriHate are TAPT from the AfriEmo dataset. DATPT + TAPT for AfriEmo results is TAPT from the AfriSenti dataset. Reported results are the Macro F1 score. Blank values (-) indicate that the specific dataset does not cover the language.
Results Table 4 shows the TAPT results across datasets and languages. As a result, all TAPT performs better than the base model and equal or competitive results with the DAPT. This indicates we can achieve better results for our targeted evaluation task using very small, quality, and task-related data. In our case, using AfriSenti data without the labels as a TAPT is very helpful for the fine-grained multi-label emotion classification task. For the AfriSenti sentiment evaluation task, TAPT with AfriHate data has better average results than TAPT with AfriEmo. For the AfriHate evaluation, TAPT in AfriSenti data has better average results than TAPT with AfriEmo. In addition to the task similarity, this improvement might be affected by the number of total instances in each dataset (the total instances of AfriSenti is 107,694, AfriEmo 70,859, and AfriHate 90,455) and the vocabulary similarity across the datasets.
6.3 Combining DAPT and TAPT
We explore the effect of both adaptation techniques by combining DAPT and TAPT. In this approach, we apply DAPT to the base model and then the TAPT training. These phases of pretraining add up to make this approach the more computationally expensive setting.
Results The results are shown in Table 4, DAPT + TAPT column. The results show that the subjective NLP tasks benefit from the combined DAPT and TAPT approaches. DAPT followed by TAPT achieves the best performance. However, first adapting the model to the domain (DAPT), then applying TAPT would be susceptible to catastrophic forgetting of the domain-relevant corpus Yogatama et al. (2019); alternate methods of combining the procedures may result in better downstream performance. This is shown from the summarized result Table 5 that sometimes the DATP + TAPT result performs less than the DAPT-only results, while it is better than the baseline results.
Dataset Models Avg.
AfriSenti AfroXLMR 51.39
   + DAPT (AfroXLMR-Social) 56.85
   + TAPT (AfriEmo) 55.72
   + TAPT (AfriHate) 55.83
   + DAPT + TAPT(AfriHate) 56.91
   + DAPT + TAPT(AfriEmo) 57.73
AfriEmo AfroXLMR 44.30
   + DAPT (AfroXLMR-Social) 51.48
   + TAPT (AfriSenti) 49.14
   + TAPT (AfriHate) 47.12
   + DAPT + TAPT(AfriSenti) 49.84
   + DAPT + TAPT(AfriHate) 48.93
AfriHate AfroXLMR 67.03
   + DAPT (AfroXLMR-Social) 70.56
   + TAPT (AfriSenti) 67.30
   + TAPT (AfriEmo) 67.73
   + DAPT + TAPT(AfriSenti) 66.55
   + DAPT + TAPT(AfriEmo) 67.18
Table 5: Summary results of Table 4, the average of the Macro F1 score across languages. Boldface values are the overall best scores for the specific dataset, and results with underlines are the best cross-TAPT dataset.
6.4 AfroXLMR-Social Vs. LLMs
This section compares our AfroXLMR-Social results with state-of-the-art open-source and proprietary Large Language Models (LLMs). Table 6 shows the summarized average results across tasks. AfroXLMR-Social leads the performance over LLMs across the three targeted tasks. Figures 2, 3, and 4 present the comparison results for the base model (AfroXLMR), AfroXLMR-Social, and both open-source and proprietary LLMs across the AfriSenti, AfriEmo, and AfriHate datasets, respectively. The result indicates that the domain-specialized encoder-only model has better or comparable results with LLMs. Generally, it means that we still need encoder-only models for low-resourced African languages and suggests that future LLMs are expected to include more training data for underrepresented languages.
Figure 2: AfriSenti Macro F1 results from AfroXLMR-Social and LLMs. The results for LLMs are based on zero-shot evaluations, selecting the best results from five different types of prompts. The benchmark results for the LLMs are taken from the AfroBench Ojo et al. (2025) leaderboard.
Figure 3: AfriEmo Macro F1 results from AfroXLMR-Social and LLMs. The results for LLMs are based on zero-shot evaluations, selecting the best results from five different types of prompts. The benchmark results for the LLMs are taken from the SemEval-2025 Task 11 dataset papers Muhammad et al. (2025b); Belay et al. (2025b).
Figure 4: AfriHate Macro F1 results from AfroXLMR-Social and LLMs. The results for LLMs are based on zero-shot evaluations, selecting the best results from five different types of prompts.. The baseline results for the LLMs are taken from the AfroBench Ojo et al. (2025) leaderboard.
Model AfriSenti AfriHate Model AfriEmo
Gemma-1.1-7B 39.7 24.3 LaBSE 35.7
Llama-2-7B 38.9 21.9 RemBERT 26.8
Llama-3-8B 41.8 27.9 XLM-R 23.4
LLaMAX3-8B 49.8 28.6 mBERT 23.0
Llama-3.1-8B 41.8 23.6 mDeBERTa 26.7
gemma-2-9B 55.5 29.9 Qwen2.5-72B 35.3
Aya-101-13B 57.1 30 Dolly-v2-12B 21.1
gemma-2-27B 58.6 45.5 Mixtral-8x7B 31.4
Llama-3.1-70B 46.9 49 Llama-3.3-70B 38.3
Gemini-1.5 pro 62.6 62.1 DeepSeek-70B 36.6
GPT-4o 62.6 63.5 AfroXLMR 44.3
AfroXLMR-Social 57.7 68.8 AfroXLMR-Social 51.5
Table 6: Summary results on fine-tuned models and LLMs. The results show the average Macro F1 score across all languages in the corresponding datasets: AfriSenti - average of 14 languages, AfriHate - average of 15 languages, and AfriEmo is the average of 17 languages.
7 Conclusion
This work explores the effects of the domain adaptive pretraining (DAPT) and task adaptive pertaining (TAPT) across three subjective tasks of African languages. We created the AfriSocial corpus, a social domain-specific corpus sourced from X and news. Using AfriSocial, we further developed the AfroXLMR-Social language model, which specialized in the social domain. We improve the performance of evaluated tasks, sentiment analysis (AfriSenti), emotion analysis (AfriEmo), and hate speech classification (AfriHate) using DAPT and TAPT techniques. We show that pretraining the model towards a small domain-specific corpus and related task-relevant data can provide significant improvements. While the combination of the two methods, DAPT + TAPT, also achieves better results than baseline models, TAPT followed by DAPT would be susceptible to catastrophic forgetting of the task-relevant corpus. We achieve better state-of-the-art results using a small domain-related corpus from the encoder-only model than state-of-the-art large-language models (LLMs). AfriSocial and AfroXLMR-Social will support the development of African languages in the NLP and improve similar-sourced tasks. It opens further domain explorations as the AfriSocial X and news domains are also available separately.
Limitations
Our work is not without limitations. We identify the following limitations with its future suggestions.
Domain Coverage.
This work explores the social media X data and news on downstream more subjective tasks such as sentiment analysis, emotion analysis, and hate speech classification. Other outside-the-domain data, such as the health domain corpus on social media-based tasks and other cross-domain evaluations, are another interesting way to explore. We have used a total of 3.5M sentences for the domain adaptation (DAPT) exploration, with a considerable variation in data statistics among languages. More balanced data from the sources across languages might provide more insights into the evaluation.
Evaluation Tasks.
In this work, our evaluation is limited to three subjective tasks: sentiment analysis, emotion analysis, and hate speech classification, as we aimed to show the effects of DATP and TAPT approaches on the social domain. This work can be extended using the proposed approaches and the AfriSocial domain-specific pre-training corpus to evaluate other similar tasks.
Appendix
Appendix A Languages Covered in Evaluation
Table 7 shows the 19 language details we evaluated in this work across different tasks.
Language ISO Subregion Spoken in Lang. family Script # Speakers
Afrikaans afr South Africa South Africa, Namibia, Botswana Indo-European Latin 7M
Amharic amh East Africa Ethiopia, Eritrea Afro-Asiatic Ethiopic 57M
Algerian Arabic arq North Africa Algeria Afro-Asiatic Arabic 36M
Moroccan Arabic ary North Africa Morocco Afro-Asiatic Arabic/Latin 29M
Hausa hau West Africa Northern Nigeria, Ghana, Cameroon Afro-Asiatic Latin 77M
Igbo ibo West Africa Southeastern Nigeria Niger-Congo Latin 31M
Kinyarwanda kin East Africa Rwanda Niger-Congo Latin 10M
Oromo orm East Africa Ethiopia Afro-Asiatic Latin 37M
Nigerian Pidgin pcm West Africa Nigeria, Ghana, Cameroon English-Creole Latin 121M
Mozambican Portug. ptMZ Southeastern Africa Mozambique Indo-European Latin 13M
Somali som East Africa Ethiopia, Kenya, Somalia Afro-Asiatic Latin 22M
Swahili swa East Africa Tanzania, Kenya, Mozambique Niger-Congo Latin >71M
Tigrinya tir East Africa Ethiopia, Eritrea Afro-Asiatic Ethiopic 9M
Twi twi West Africa Ghana Niger-Congo Latin 9M
Makhuwa vmw East African Mozambique, Tanzania Niger–Congo Latin 7M
Xitsonga tso Southern Africa South Africa, Zimbabwe, Mozambique Niger-Congo Latin 7M
Xhosa xho Southern Africa South Africa, Zimbabwe, Lesotho Niger-Congo Latin 19M
Yoruba yor West Africa Southwestern, Central Nigeria, Togo Niger-Congo Latin 46M
Zulu zul Southern Africa South Africa Niger-Congo Latin 29M
Table 7: Additional information on the African languages evaluated in this work: ISO-3 digit language code, region spoken, the family of the language, its script, and number of L1 and L2 speakers.
Appendix B Evaluation Data statistics
Table 8 shows the train-test split of the evaluation datasets AfriSenti, AfriEmo, and AfriHate.
Lang. AfriSenti AfriEmo AfriHate
Train Dev Test Total Train Dev Test Total Train Dev Test Total
afr - - - - 2107 98 1065 3270 - - - -
amh 5985 1498 2000 9483 3549 592 1774 5915 3467 744 747 4958
arq 1952 415 959 3062 901 100 902 1903 716 211 323 1250
ary 5584 1216 2962 9762 1608 267 812 2687 3240 695 699 4634
hau 14173 2678 5304 22155 2145 356 1080 3581 4566 1029 1049 6644
ibo 10193 1842 3683 15718 2880 479 1444 4803 3419 774 821 5014
kin 3303 828 1027 5158 2451 407 1231 4089 3302 706 714 4722
orm - 397 2097 2494 3442 575 1721 5738 3517 763 759 5039
pcm 5122 1282 4155 10559 3728 620 1870 6218 7416 1590 1593 10599
ptMZ 3064 768 3663 7495 1546 257 776 2579 - - - -
som - - - - 3392 566 1696 5654 3174 741 745 4660
swa 1811 454 749 3014 3307 551 1656 5514 14760 3164 3168 21092
tir - 399 2001 2400 3681 614 1840 6135 3547 760 765 5072
twi - - - - - - - - 2564 639 698 3901
vmw - - - - - 1551 258 777 2586 - - - -
xho 805 204 255 1264 - 682 1594 2276 2502 559 622 3683
yor 8523 2091 4516 15130 2992 497 1500 4989 3336 724 819 4879
zul - - - - - 875 2047 2922 2940 640 728 4308
Table 8: Dataset distribution across different languages - AfriEmotion (train, dev, test, and total) and AfriSenti dataset. We adopt the same train-test-dev split from the data source papers: AfriEmo Muhammad et al. (2025c), AfriSenti Muhammad et al. (2023), and AfriHate Muhammad et al. (2025a).
Appendix C Baseline results from encoder-only LMs
We evaluate various encoder-only language models with the more difficult multi-label emotion classification task to select the best encoder-only model for continual learning. We found that AfroXLMR is better for the low-resourced African languages, and we continue our DAPT and TAPT training settings from this model. Table 9 shows baseline results for the AfriEmo dataset from different multilingual encoder-only models.
Model afr amh ary hau ibo kin orm pcm ptMZ som swa tir vmw yor xho zul
Baselines from general Multilingual models
LaBSE 37.76 63.72 45.81 58.49 45.90 50.64 43.30 51.30 36.95 41.82 27.53 47.23 21.13 32.55 31.39 18.16
RemBERT 37.14 63.83 47.16 59.55 47.90 46.29 12.63 55.50 45.91 45.93 22.65 46.28 12.14 9.22 12.73 15.26
mBERT 26.87 26.69 36.87 47.33 37.23 35.61 39.84 48.42 14.81 31.13 22.99 25.16 10.28 21.03 17.08 13.04
mDeBERTa 16.66 44.22 38.00 48.59 31.92 38.00 28.48 46.21 21.89 34.91 22.84 30.35 11.13 17.88 22.86 13.87
Baselines from African language-centric models
AfroLM 21.60 54.78 30.35 57.31 42.46 38.97 41.84 47.12 17.81 32.43 20.08 38.22 15.98 24.31 13.67 10.72
AfriBERTa 22.90 60.05 30.85 61.09 49.05 46.35 53.69 50.29 23.15 44.92 24.36 49.00 20.29 34.52 13.86 8.50
AfroXLMR 43.66 68.97 47.62 64.30 26.27 52.39 52.28 55.39 22.09 48.78 30.74 57.22 21.18 28.65 13.52 6.90
Continual pretraining from XLM-RoBERa-Large
XLMR-L 38.69 54.99 38.31 52.99 38.72 35.06 26.67 53.77 9.29 39.95 6.62 18.58 13.53 2.79 7.90 9.27
+ DAPT 25.94 60.08 44.41 59.81 42.91 44.61 43.61 53.74 30.57 41.69 27.83 49.55 8.44 18.05 4.75 9.27
+ TAPT 39.65 62.23 47.61 63.20 47.56 39.09 45.25 57.78 36.68 38.68 30.28 44.74 14.89 26.32 10.03 12.96
+ DAPT +TAPT 23.00 60.81 42.91 60.44 43.44 41.53 45.26 53.90 29.83 40.05 24.02 48.28 9.32 18.53 6.19 4.63
Continual Fine-tuning from AfroXLMR-large
AfroXLMR 43.66 68.97 47.62 64.30 26.27 52.39 52.28 55.39 22.09 48.78 30.74 57.22 21.18 28.65 13.52 6.90
+ DAPT 44.57 71.67 52.63 70.74 54.54 56.73 61.38 59.93 36.80 54.86 34.35 60.71 22.08 39.26 8.54 6.72
+ TAPT 49.61 69.56 50.21 66.93 53.13 53.27 56.43 56.71 37.20 50.33 33.02 55.72 21.73 34.34 4.58 13.15
+ DAPT + TAPT 44.17 71.04 51.63 69.77 54.26 54.47 58.75 58.89 37.76 52.65 32.74 57.12 19.80 35.89 21.30 13.81
Table 9: AfriEmo detail results using AfriSocial as DAPT and AfriSenti as TAPT. xho and zul languages have no training set, and the results are in zero-shot. The results are from the average scores of five runs.
Appendix D Training Details of DAFT and TAPT
Hyper-parameters for adapters We trained the task adapter using the following hyper-parameters: batch size of 32, 10 epochs, and learning rate of 5e-5. For TAPT, the parameters are similar to those of DAPT, except that the batch size is 8. We used their tokenizer, for XLMR - XLMR tokenizer, and AfroXLMR - AfroXLMR tokenizer. PyTorch was used for fine-tuning, and pre-trained models were sourced from Hugging Face. The domain adaptive fine-tuning training is trained on three distributed-based GPUs. Following standard practice, we pass the final layer [CLS] token representation to a task-specific feedforward layer for prediction with three epochs. The reported results from fine-tuned pre-trained language models are the average results of five runs.
Appendix E Large Language Model Details
Multilingual Encoder-only, open-source, and proprietary model names and their sources are mentioned below. The results from LLMs are used from the work Ojo et al. (2025); Muhammad et al. (2025b). All open-source LLMs are instruction-tuned versions. The various evaluation prompts are presented in the original works mentioned above.
E.1 Encoder-only Langauge Models
  • LaBSE Feng et al. (2022) - sentence-transformers/LaBSE
  • RemBERT Chung et al. (2020) - google/rembert
  • XLM-RoBERTa Conneau et al. (2020b) - FacebookAI/xlm-roberta-base (large)
  • mDeBERTa He et al. (2021) - microsoft/mdeberta-v3-base
  • mBERT Libovický et al. (2019)- google-bert_bert-base-multilingual-cased
  • AfriBERTa Ogueji et al. (2021) - castorini/afriberta_large
  • AfroXLMR Alabi et al. (2022) - Davlan/afro-xlmr-large-61L (76L)
  • AfroLM Dossou et al. (2022)- bonadossou/afrolm_active_learning
E.2 Open-source LLMs
  • Aya-101 13B Üstün et al. (2024) - CohereLabs/aya-101
  • Llama 2 7B Chat Touvron et al. (2023) - meta-llama/Llama-2-7b-chat-hf
  • Llama 3 8B Grattafiori et al. (2024) - meta-llama/Meta-Llama-3-8B-Instruct
  • Llama 3.1 (8B, 70B) Grattafiori et al. (2024) - meta-llama/Llama-3.1-8B-Instruct and meta-llama/Llama-3.1-70B-Instruct, respectively.
  • Gemma 1.1 7B Mesnard et al. (2024) - google/gemma-1.1-7b-it
  • Gemma 2 (9B, 27B) Team et al. (2024) - google/gemma-2-2b-it and google/gemma-2-27b-it
  • DeepSeek-R1-70 Guo et al. (2025) - deepseek-ai/DeepSeek-R1-Distill-Llama-70B
  • Mistral-8x7B Jiang et al. (2024) - mistralai/Mixtral-8x7B-Instruct-v0.1
  • Qwen2.5-72B Yang et al. (2024) - Qwen/Qwen2.5-72B-Instruct
  • Dolly-v2-12B Conover et al. (2023) - databricks/dolly-v2-12b
E.3 Propritary LLMs
  • Gemini 1.5 Pro Georgiev et al. (2024) - Gemini 1.5 Pro 002 accessed via Google API
  • GPT-4o (Aug) Hurst et al. (2024) - the August 2024 version of the model is accessed through the OpenAI API
Appendix F AfriSenti detail results
Table 10 shows all sentiment analysis results (AfriSenti dataset).
Models Languages Avg.
amh arq ary hau ibo kin orm pcm por swa tir tso twi yor
Fine-tuned encoder-only models from the AfroXLMR baseline
AfroXLMR 50.1 52.2 52.9 79.3 76.9 71.0 20.1 50.5 60.9 28.3 22.5 35.4 47.2 72.3 51.4
AfroXLMR-Social 57.2 64.6 62.3 62.7 79.8 72.7 24.3 52.1 64.8 61.4 24.5 38.8 56.0 74.6 56.9
TAPT-Emo 54.9 62.4 64.1 80.7 80.0 71.5 23.5 51.0 64.1 59.3 10.8 37.9 47.7 72.2 55.7
TAPT-Hate 54.5 59.4 52.8 80.1 78.3 69.7 43.0 50.3 62.8 57.3 16.2 36.2 50.2 70.9 55.8
DAPT+TAPT-Emo 55.8 63.4 63.1 82.7 80.4 69.5 28.9 52.0 63.8 54.9 28.9 36.7 54.5 73.7 57.7
DAPT+TAPT-Hate 56.3 59.7 62.1 82.0 80.2 70.1 23.6 51.2 62.1 58.5 21.9 40.0 55.8 73.4 56.9
Prompt-based proprietary models
LLaMAX3-8B 55.2 55.5 51.0 61.7 54.6 53.2 33.6 56.0 41.3 54.1 43.5 48.0 39.0 50.4 49.8
Llama-2-7B 25.5 44.9 44.0 38.2 33.6 35.4 24.7 60.8 31.2 33.8 33.5 46.1 48.9 43.7 38.9
Llama-3.1-70B 40.0 47.5 53.5 52.6 52.2 48.5 41.4 52.6 35.9 61.5 28.2 43.3 45.8 54.3 47.0
Llama-3.1-8B 66.4 57.1 51.9 55.4 50.1 48.7 35.9 64.2 33.6 54.3 49.8 48.8 42.3 50.9 50.7
Llama-3-8B 46.3 51.0 46.1 38.5 36.1 38.4 28.2 60.2 27.9 37.8 38.0 43.3 47.7 45.1 41.8
Aya-101 76.8 67.8 58.1 61.2 47.5 61.1 37.4 70.1 48.8 47.5 71.2 50.8 44.7 57.0 57.1
Gemma-1.1-7B 24.4 43.1 42.0 37.9 34.7 32.0 25.9 66.5 37.4 37.0 32.4 50.0 48.7 43.8 39.7
Gemma-2-27B 70.7 65.8 59.0 64.8 60.4 59.1 37.3 76.0 42.8 55.6 58.9 50.0 54.3 65.5 58.6
Gemma-2-9B 70.1 62.0 56.4 61.4 58.2 56.1 37.9 66.8 46.6 58.7 55.4 43.7 48.1 55.4 55.5
Prompt-based proprietary models
Gemini-1.5 77.5 70.9 63.7 70.1 56.9 68.3 42.8 74.5 46.4 55.2 70.2 55.9 49.3 74.3 62.6
GPT-4o 75.6 72.3 61.2 68.6 67.8 71.6 43.1 67.1 62.1 57.9 61.5 46.5 51.3 70.2 62.6
Table 10: AfriSenti Model Performance Across Various Languages
Appendix G AfriHate results
Table 11 shows all hate speech classification results (AfriHate dataset).
Models Languages Avg.
amh arq ary hau ibo kin orm pcm som swa tir twi xho yor zul
Fine-tuned encoder-only models from the AfroXLMR baseline
AfroXLMR 73.5 43.4 75.1 81.6 82.8 75.3 67.2 64.9 55.7 91.5 50.2 46.9 50.9 53.4 54.5 64.5
DAPT 78.6 46.0 75.6 80.8 88.1 78.8 74.1 67.6 55.6 91.2 55.9 48.4 59.2 77.9 55.4 68.9
TAPT-Emo 73.1 46.2 75.1 77.8 87.7 77.5 69.0 67.7 57.3 91.9 54.2 49.3 55.1 53.7 56.0 66.1
TAPT-Senti 73.3 43.8 71.7 81.9 86.5 76.3 67.7 66.9 55.0 91.3 57.0 48.9 51.3 54.5 55.0 65.4
DAPT + TAPT-Emo 78.1 44.2 77.5 82.1 87.7 78.4 71.6 69.9 56.8 91.2 32.7 49.0 58.7 54.8 55.5 65.9
DAPT + TAPT-Senti 77.6 43.8 76.1 81.6 79.4 77.9 72.0 67.5 53.7 91.5 41.2 48.2 54.6 54.8 55.8 65.0
Prompt-based proprietary models
Gemma-1.1-7B 23.0 27.4 24.5 26.0 16.7 29.9 27.9 30.2 27.2 27.4 17.3 14.2 23.3 25.0 22.5 24.3
Llama-2-7B 14.5 22.4 22.2 24.4 20.2 22.4 31.3 9.4 27.1 24.8 11.7 15.8 24.8 23.1 26.8 21.9
Llama-3-8B 26.5 31.8 28.5 24.5 19.7 36.5 37.1 38.8 17.8 34.3 28.4 14.4 25.0 25.9 28.4 27.9
LLaMAX3-8B 37.2 33.6 31.5 30.7 19.4 38.2 38.2 34.4 27.6 28.9 27.4 13.9 23.7 24.4 29.0 28.6
Llama-3.1-8B 23.3 30.7 22.9 25.4 13.9 31.9 35.7 24.9 26.7 21.7 21.9 9.9 22.4 19.4 23.3 23.6
gemma-2-9b 33.2 33.8 33.2 24.1 25.1 33.6 26.7 54.9 13.6 46.4 26.8 29.1 20.0 30.5 20.1 29.9
Aya-101-13B 31.3 32.1 28.9 33.3 22.1 32.8 26.8 37.8 35.8 41.3 29.6 13.8 28.7 26.8 29.8 30.0
Gemma-2-27B 48.4 49.1 53.8 34.8 42.8 52.7 39.8 60.9 39.6 70.9 35.4 38.1 30.6 54.0 35.0 45.5
Llama-3.1-70B 53.0 57.0 60.6 41.2 48.4 50.9 44.6 62.4 39.8 67.0 41.0 37.9 32.7 56.2 46.3 49.0
Prompt-based proprietary models
Gemini-1.5 pro 56.1 70.6 68.2 61.4 66.9 64.2 57.6 65.0 60.8 80.5 37.5 50.6 58.0 73.1 55.4 62.1
GPT-4o 56.0 69.7 75.5 59.2 69.7 60.1 53.5 65.2 68.5 78.0 42.4 51.2 63.7 74.5 58.7 63.5
Table 11: AfriHate Model Performance Across Various Languages
Appendix H AfriEmo detail results
Table 12 shows all fine-grained multi-label emotion classification results (AfriEmo dataset).
Models Languages Avg.
afr amh arq ary hau ibo kin orm pcm ptMZ som swa tir vmw yor xho zul
Fine-tuned encoder-only models from the AfroXLMR baseline and others
AfroXLMR 43.7 69.0 44.9 47.6 64.3 26.3 52.4 52.3 55.4 22.1 48.8 30.7 57.2 21.2 28.7 13.5 6.9 40.3
+ DAPT 44.6 71.7 51.3 52.6 70.7 54.5 56.7 61.4 59.9 36.8 54.9 34.4 60.7 22.1 39.9 8.5 6.7 46.3
+ TAPT-Senti 49.6 69.6 49.0 50.2 66.9 53.1 53.3 56.4 56.7 37.2 50.3 33.0 55.7 21.7 34.3 14.6 13.2 45.0
+ TAPT-Hate 47.2 67.2 48.4 48.8 61.2 51.3 53.5 52.2 56.6 31.0 49.6 31.9 55.8 19.3 32.9 11.0 9.9 42.8
+ DAPT + TAPT-Senti 44.2 71.0 48.7 51.6 69.8 54.3 54.5 58.8 58.9 37.8 52.7 32.7 57.1 19.8 35.9 21.3 13.8 46.1
+ DAPT + TAPT-hate 46.2 70.8 46.6 48.8 70.0 54.2 53.5 56.8 58.2 34.8 52.7 31.4 57.1 19.7 33.0 9.4 5.0 44.0
LaBSE 35.1 63.7 35.9 42.8 38.5 18.1 30.4 43.3 33.3 31.4 41.8 21.7 47.2 9.7 11.6 31.4 18.2 32.6
RemBERT 35.0 63.8 33.8 35.5 32.0 7.5 18.4 12.6 1.0 29.7 45.9 19.0 46.3 5.2 5.3 12.7 15.3 24.7
XLM-R 41.7 46.9 35.9 33.9 16.7 10.4 13.1 17.9 21.1 7.3 25.4 16.9 35.9 12.7 6.6 11.5 10.9 21.5
mBERT 17.0 26.7 31.4 24.8 15.6 9.9 20.9 39.8 22.6 13.5 31.1 18.6 25.2 12.1 9.6 17.1 13.0 20.5
mDeBERTa 33.3 44.2 35.9 36.3 32.8 9.5 17.3 28.5 25.4 24.5 34.9 14.9 30.4 11.7 10.0 22.9 13.9 25.1
Prompt-based proprietary models
Qwen2.5-72B 60.2 37.8 37.8 52.8 43.8 37.4 32.0 31.6 38.7 40.4 28.6 27.4 31.1 20.4 25.0 29.6 22.0 35.1
Dolly-v2-12B 23.6 5.1 38.6 24.3 29.4 24.3 19.7 22.9 34.4 16.7 19.8 17.6 1.5 16.0 16.0 24.1 14.7 20.5
Llama-3.3-70B 61.3 42.8 55.8 45.0 50.9 33.2 34.4 29.8 48.7 34.1 32.5 29.5 32.9 19.0 23.7 30.8 21.5 36.8
Mixtral-8x7B 53.7 29.0 45.3 35.1 40.4 31.9 26.4 24.3 45.6 36.5 25.6 26.5 27.2 19.0 19.7 22.9 20.4 31.1
DeepSeek-R1-70B 43.7 36.9 50.9 47.2 51.9 32.9 32.5 28.2 45.0 39.6 26.6 33.3 26.5 19.1 27.4 29.1 20.4 34.8
Table 12: AfriEmo Model Performance Across Various Languages
Appendix I AfriSocial Data Sources
X (Twitter) Source
There is an X domain corpus and model for high-resource languages such as XLM-T Barbieri et al. (2022) to evaluate and improve task datasets sourced from X. However, there is a scarcity of corpora specializing in the social domain for low-resource African languages. The tweets are scraped over a different time until June 2023, before the change of an X policy that restricts their data for academic research.
News Source
News platforms are the most common data source for African languages. Companies also stream their news on the X platform. While more formal, news websites also provide a platform for public discourse, comments, and reactions, often including opinion pieces and user-generated comments. The source news websites are British Broadcasting Corporation (BBC) news [^3] https://www.bbc.com/, Akan news [^4] https://akannews.com, Global Voice News [^5] https://mg.globalvoices.org/, isolezwelesixhosa [^6] https://www.isolezwelesixhosa.co.za/, isolezwe [^7] https://www.isolezwe.co.za/, and others.